X (t) = xd_ya y (t) = dy-xa y=- x=1 Find 2 x X ㅓ >
Answer #1

and こ XCt) = xya Y(0) = -1 Y(H) = 2x-x² 7. X(0)=1 X(0) = (x(0)) ? (X(01) 1-4)= 0 = x(0) = Y0) - 2900) – (X() 261) - (119